OpenStudy (anonymous):

Intially there are 100 bacteria. The bacteria doubles every 90 minutes. Determine the constant r in f(t)=Ce^(rt)

OpenStudy (anonymous):

C=100 \[200=100*e ^{(r*90)}\]\[\frac{ 200 }{ 100 }=e ^{(r*90)}\]\[\ln 2=r*90\]\[r=\frac{ \ln 2 }{ 90 }\]